# Interpretable Multi-Source Data Fusion Through Latent Variable Gaussian Process

### v1

@@ 4.4.4 Enchancing Modeling Capabilities Through Targeted Source Selection Based on Latent Variables
It is generally acklowledged in the ML community that the prediction performance of the models improve when more data is available. Thus, there has been great efforts of data collection and curation in the ML community. Although this notion of more the data better the model is generally true, there could be cases where more data can actually hurt the ML models and result in poor prediction results. This phenomenan can be especially true for ML applications in materials science where the fundamentals of the generated data, specifically for experimental data, are usually either unknown to the model builders or not disclosed by the data generators due to many reasons such as errors, confidentiality and so on. Therefore, a major challenge still exists regarding the quality, and applicability of the available data. Thus, during the ML model building process, researchers are more inclined to trust and use the data they either generate themselves or they know the fundametals of. However, this comes with two challenges. First, generating materials data whether experimental or simulation can be expensive and time consuming, which could result in poor ML models further down the line due to low amounts of available training samples. As a result, researchers also rely on other data sources to build their models, in addition to using the data they generate themselves. However, as mentioned previously, it is not always plausible to know the quality and the fundamentals of the data generation process, meaning that there could be notable differences and discrepencies in the data sources which could significantly influence the performance of the ML models.
@@
So far, it has been demonstrated that the LVGP modeling brings numerous advantages to multi-source data fusion compared to some of the alternative approaches. One of the key advantages of the LVGP comes from the physically interpretable latent spaces that can possibly explain the relationships and differences between different sources of information. As a result, it is proposed that the latent space provided by the LVGP model can act as both a validation and an anomaly detection mechanism to remove the data samples from the sources that could potentially harm the ML model that will be used for further optimization tasks on the source of information that one has control over. For the data sources available in the FeCrAl alloy study, the authors generated and have further control over only on the GE Data source. Therefore, the further material design optimization tasks can only take place using the experimental procedures available at GE Research. However, to improve the quality of the ML modelling, numerous data sources have also been collected and used for LVGP model building. Looking at the latent spaces from LVGP model in Figure 7a, it is observed that some data sources, Yang, Kobayashi, and Pinkas to be specific, are located at a much further point (higher D values) compared to the other sources when GE Data is selected as the reference point. As discussed before, this result can indicate that the experimental procedures taken place by those sources can be different than the ones that are implemented for GE Data. However, this does not necessarily state that there are fundamental errors or data quality issues regarding the three data sources but could potentially indicate that the sources closer to the GE Data can contribute to model accuracy whereas others can harm the ML model for predictions on GE Data. Therefore, to investigate the influnece of targeted source selection on the seperate testing GE Data set, the three sources that are further from the GE Data (Yang, Kobayashi, and Pinkas), is removed from the training sets and the LVGP model is re-trained. For ease of notation, the new LVGP model is denoted as LVGP-Target. Table 4 shows the comparision between the prediction accuracy made by the three models, LVGP, GP-GE, and LVGP-Target on the seperate 36 GE Data testing set. Although both LVGP and GP-GE have already demonstrated remarkable accuracy results on the GE Data, LVGP-Target provides notable improvement in accuracy, specifically reducing NRMSE by almost 20% and 7% in comparision to LVGP and GP-GE models, respectively. The parity plot of the predictions made by the LVGP-Target model are shown in Supplementary Figure 2. Considering that the evaluation of properties through experimenation can be relatively expensive, the notable increase in accuracy provided by the targeted source selection through LVGP can provide notable cost and time savings for further downstream tasks. Overall, in addition to providing accurate data fusion modeling, LVGP can be employed as a validation and data filtering mechanism to improve multi-source modeling.
@@ 5.1 Overview
Electric and hybrid electric transportion has been garnering a lot of attention in recent times. One of the key focueses has been developing more denser and powerful motors to provide the propulsion capacity needed. To increase the power density and capacity of the motor, the housed permanent magnet plays an essential role. With the direct increase in magnetic characteristics of fundamental magnetic material, high operational envelope of motors can be achieved. An overview of performance of magnetic alloy systems is given through [42]. The work identifies alloys that have higher magnetic properties measured through saturation magnetization and that have higher operating temperatures measured through Curie temperature. From the array of magnetic materials, it was be observed SmCo (Samarium-Cobalt) alloy, also referred as SmCoFe, exhibit relatively high stability at high temperatures which is essential to withstand the hot and harsh environments of aircraft engines. Hence the research community have favored the exploration of SmCoFe systems for aerospace application.
@@ 5.2 Data Description
As with the previous materials case study, the first step is to collect data from open research literature such as papers, patents, and other sources of information that are restricted to SmCoFe alloy system. From an extensive search, three primary sources of information were identified. Unfortunately, the in-house data testing for SmCoFe systems are not readily available but currently on-going. Table 5 summarizes the sources of data. A total of 71 data points were collected. One distinction from the previous case study is that the data from the one of the patents is further differentiated into two datasources (Patent T2 and Patent T4). The differentiation was necessitated by additional changes to the post-processing parameters that were done.
@@
Name of Data Source
Number of Data Samples
Source of Data
Patent T2
$20$
Table 2 from Fujiwara, Teruhiko, Hiroaki Machida, and Hideyuki Yoshikawa. "Rare earth-cobalt permanent magnet." U.S. Patent Application No. 14/643,875. [43]
Patent T4
$21$
Table 4 from Fujiwara, Teruhiko, Hiroaki Machida, and Hideyuki Yoshikawa. "Rare earth-cobalt permanent magnet." U.S. Patent Application No. 14/643,875. [43]
Prop of SmCeCoFeCu
$30$
Senno, Harufumi, and Yoshio, Tawara. "Permanent-magnet properties of Sm-Ce-Co-Fe-Cu alloys with compositions between 1-5 and 2-17." IEEE Transactions on Magnetics 10.2 (1974): 313-317. [44]
Total
$71$
@@ 5.3 Model Development
Simlar to the FeCrAl model development approach, GP and LVGP models are trained on the collected data. The common inputs for the GP and LVGP models are the composition of the alloy constituents defined as the weight percentages (Samarium, Zirconium, Copper, Iron, Cerium, Cobalt) and the testing temperature of the alloy, leading to seven (7) unique input parameters. The energy density and power capacity of the magnet are experimentally measured and quantified through energy product (EP). As with the previous case study, the inputs for the LVGP contain an additional qualitative input characterizing each source of data as distinct categorical level along all the other 7 inputs used for the GP model. The outputs for both the models remain the same.
@@ 5.4 Results and Discussion
Two seperate studies are conducted as previously done with FeCrAl study. Within the first study, generazability and the prediction capability of the models through one source that could be controlled are investigate. In this case, altough there is no in-house experiments that could be controlled, Prop of SmCeCoFeCu is selected as in-house experiments (reference information source) as it is the source with the highest number of available samples. First, generazability of GP and LVGP is compared through 5-fold cross validation using the 65 samples out of 71 available samples, and tested the models over a 6 sample seperate testing set that are selected from the Prop of SmCeCoFeCu source. For CV tasks in LVGP, stratified sampling is used to make sure datasets used for training and testing represent the overall distribution of the information sources. The training dataset and testing dataset remain identical for both models. The results of this study are highlighted in Section 5.4.1.
@@
In the second study, seperate single-source GP model built using only the Prop of SmCeCoFeCu source and compared with the multi-source LVGP data fusion model. The performance of the models were compared according to the same 6 seperate testing set. For ease of notation, this single source GP model is named as GP-Prop. The results of this study are demonstrated in Section 5.4.2.
5.4.1 Multi-Source Data Fusion Comparision: LVGP and GP
@@
Table 6 summarize the performance of the GP and LVGP model on the SmCoFe alloy dataset. From the comparison of the NRMSE across the training and cross validation datasets, it can be observed that the multi-source LVGP model shows improved performance over the GP model. Therefore, it can be concluded that explicitly including the source of data could enable accurate, efficient and generazable models. Looking at the testing NRMSE metric, it is observed that GP model performs slightly better compared to the LVGP model. The parity plots from the CV study are also shown in Supplementary Figure 3. Moreover, from 45 degree parity plots, it is observed that both models make similar predictions on both the testing and training set (Panels A and B of Figure 8). On the other hand, it can clearly observed that LVGP provides lower prediction uncertainities as the predictions have lower confidence bounds. Once again, this could be highly advantegeous for further downstream materials design optimization studies as the information from other sources can provide a more informed and a confident ML model.
@@
Table 6: Predictive performance comparision for GP and LVGP on SmCoFe Alloys
Model   Type
Mean Training NRMSE
Mean CV NRMSE
Testing NRMSE
GP
$0.100$
$0.134$
$0.160$
LVGP
$0.102$
$0.128$
$0.165$
GP-Prop
$0.326$
@@
Figure 8: Predictive performance comparision between (a) GP, (b) LVGP, and (c) GP-Prop models. First row shows the parity plots of the models with respect to their respective training and Prop SmCeCoFeCu testing sets.The x and y axis show the true and predicted energy potential (EP), respectively, and error bars demonstrate the prediction uncertainity associated with each prediction. Second, third and fourth row show the Mean, Standard Deviation and 3D Response Surfaces of the models at 23%<Sm<26%, 55%<Co<63%, 1.0%<Zr<2.0%, Cu = 8.0%, Fe = 19.0%, Hf = 0%, Ce = 0% and Temperature = ${}^{\circ}$, respectively
@@
Extending the investigation of the models, the response surface from both the models are plotted across the same input space of the Prop SmCeCoFeCu source through Panels A-B in Figure 8. The response surfaces (with Prop SmCeCoFeCu as reference due to details and nuances provided in the source) of the models are smooth and well-behaved. The well-behaved nature of the response surface can be attributed to the low dissimilarity measure of each source given through Figure 9. This could indicate that all the sources of information possess high similarities. It is highlighted that since the experiments for Patent T2 and Patent T4 have been provided through the same Patent [43], it is plasuable to assume that the underlying unknown physical experimental parameters are the same for both sources of information and that LVGP was able to identify this explicitly. Furthermore, in addition to providing better predictions, a case is still made for the need of multi-source data fusion through LVGP. Within the highlighted compositional space of the response surface, the location of maximum energy product from GP and LVGP are different. For the GP model, the maximum energy product is around 17.75 MGOe with a prediction uncertainty of around 2.5 MGOe, approximately at Co(wt%) = 60. For LVGP models centered at Prop SmCeCoFeCu, the maximum energy product is around 18.00 MGOe with an uncertainty of 2.25 MGOe at approximately Co(wt%) = 63. This nuanced difference is seemingly tied to the source of the data as datasets and models are identical except for inclusion of data source. Although the mean predictions provided by the LVGP is slightly higher compared to the GP model, the predictions come with much lower uncertainity. Furthermore, for the cases where the designer has access and control over one source of information for further design optimization, the optimimum candidates provided by the GP model can be misleading and can misguide the researchers towards wrong optimum designs since there is no way to explicitly differentiate the information sources. Therefore, the reference source selection capability present in the LVGP plays an essential role in plausibly guaranteeing the optimized behavior as it is tied to the most informative source of data (identical procedures can be followed). Additionally, within the compositional space there is a difference of 3% in Co (wt%) for the maximum EP property between both models. This shift in the composition of a key rare earth element can have a significant impact in realizing such a superior alloy and downstream cost and supply chain setup.
@@
Figure 9: Latent Space of SmCoFe from LVGP. Prop SmCeCoFeCu is denoted as the reference source (*) for dissimilarity metric, $D$, calculations
5.4.2 Comparision of Multi-Source LVGP with Single-Source GP-Prop
@@
To overcome the challenge of performing optimization in the unknown information source space, one can argue to build GP models on each information source individually. To demosntrate the effectiveness of multi-source LVGP, it is compared against a single-source GP model that is built on the Prop SmCeCoFeCu source only. The single-source GP model, which is previously denoted as GP-Prop, uses the entire available training data from the Prop SmCeCoFeCu source (24 samples) for model building, and later tested on the separate 6 samples obtained from the Prop SmCeCoFeCu once again. The prediction metrics and the parity plots on testing data are shown in Table 6 (Model Types: LVGP and GP-Prop) and 9 b&c, respectively. From the NRMSE and parity plot results, a clear advantage of using LVGP is observed as it provides much better predictions with lower prediction uncertainities. Furthermore, looking at the response surfaces of both models, Panels B and C of Figure 8, a tone stark difference is seen. At Co (wt%)>61 (approximately), the GP-Prop model reverts to mean predictions with high prediction uncertainty, which is a common behavior observed in GP models when trying to obtain prediction for out-of-bounds inputs in the training data. However, when different sources of information are incorporated with non-overlapping input space through LVGP, the model is able to perform extrapolation with regard to the reference source. This phenomenon could be a crucial differentiator when designing next generation materials with restrictions on data-availability driven by manufacturing and experimental testing capabilities.

### v3

@@ 4.4.4 Enhancing Modeling Capabilities Through Targeted Source Selection Based On Latent Variables
It is generally acknowledged in the ML community that the prediction performances of the models improve when more data is available. Thus, there has been great efforts of data collection and curation in the ML community. Although this notion of more the data better the model is generally true, there could be cases where more data can actually hurt the ML models and result in poor prediction results. This phenomenon can be especially true for ML applications in materials science where the fundamentals of the generated data, specifically for experimental data, are usually either unknown to the model builders or not disclosed by the data generators due to many reasons such as errors, confidentiality and so on. Therefore, a major challenge still exists regarding the quality, and applicability of the available data. Thus, during the ML model building process, researchers are more inclined to trust and use the data they either generate themselves or they know the fundamentals of. However, this comes with two challenges. First, generating materials data whether experimental or simulation can be expensive and time consuming, which could result in poor ML models further down the line due to low amounts of available training samples. As a result, researchers also rely on other data sources to build their models, in addition to using the data they generate themselves. However, as mentioned previously, it is not always plausible to know the quality and the fundamentals of the data generation process, meaning that there could be notable differences and discrepancies in the data sources which could significantly influence the performance of the ML models.
@@
So far, it has been demonstrated that the LVGP modeling brings numerous advantages to multi-source data fusion compared to some of the alternative approaches. One of the key advantages of the LVGP comes from the physically interpretable latent spaces that can possibly explain the relationships and differences between different sources of information. As a result, it is proposed that the latent space provided by the LVGP model can act as both a validation and an anomaly detection mechanism to remove the data samples from the sources that could potentially harm the ML model that will be used for further optimization tasks on the source of information that one has control over. For the data sources available in the FeCrAl alloy study, the authors generated and only have further control over the GE Data source. Therefore, the further material design optimization tasks can only take place using the experimental procedures available at GE Research. However, to improve the quality of the ML modelling, numerous data sources have also been collected and used for LVGP model building. Looking at the latent spaces from LVGP model in Figure 8a, it is observed that some data sources, Yang, Kobayashi, and Pinkas to be specific, are located at a much further point (higher D values) compared to the other sources when GE Data is selected as the reference point. The stark differences are also observed in the source-specific response surfaces learned and provided by the LVGP (Figure 7). As discussed before, this result can indicate that the experimental procedures taken place by those sources can be different than the ones that are implemented for GE Data. However, this does not necessarily state that there are fundamental errors or data quality issues regarding the three data sources but could potentially indicate that the sources closer to the GE Data can contribute to model accuracy whereas others can harm the ML model for predictions on GE Data. Therefore, to investigate the influence of targeted source selection on the separate testing GE Data set, the three sources that are further from the GE Data (Yang, Kobayashi, and Pinkas), is removed from the training sets and the LVGP model is re-trained. For ease of notation, the new LVGP model is denoted as LVGP-Target. Table 4 shows the comparison between the prediction accuracy made by the three models, LVGP, GP-GE, and LVGP-Target on the separate 36 GE Data testing set. Although both LVGP and GP-GE have already demonstrated remarkable accuracy results on the GE Data, LVGP-Target provides notable improvement in accuracy, specifically reducing NRMSE by almost 20% and 7% in comparison to LVGP and GP-GE models, respectively. The parity plot of the predictions made by the LVGP-Target model are shown in Supplementary Figure 2. Considering that the evaluation of properties through experimentation can be relatively expensive, the notable increase in accuracy provided by the targeted source selection through LVGP can provide significant cost and time savings for further downstream tasks. Overall, in addition to providing accurate data fusion modeling, LVGP can be employed as a validation and data filtering mechanism to improve multi-source modeling.
@@ 5.1 Overview
Electric and hybrid electric transportation has been garnering a lot of attention in recent times. One of the key focuses has been developing more denser and powerful motors to provide the propulsion capacity needed. To increase the power density and capacity of the motor, the housed permanent magnet plays an essential role. With the direct increase in magnetic characteristics of fundamental magnetic material, high operational envelope of motors can be achieved. An overview of performance of magnetic alloy systems is given through [cui2018current]. The work identifies alloys that have higher magnetic properties measured through saturation magnetization and that have higher operating temperatures measured through Curie temperature. From the array of magnetic materials, it was observed that SmCo (Samarium-Cobalt) alloy, also referred as SmCoFe, exhibit relatively high stability at high temperatures which is essential to withstand the hot and harsh environments of aircraft engines. Hence the research community has favored the exploration of SmCoFe systems for aerospace application.
@@ 5.2 Data Description
As with the previous materials case study, the first step is to collect data from open research literature such as papers, patents, and other sources of information that are restricted to SmCoFe alloy system. From an extensive search, three primary sources of information were identified. Unfortunately, the in-house data testing for SmCoFe systems is not readily available but currently on-going. Table 5 summarizes the sources of data. A total of 71 data points were collected. One distinction from the previous case study is that the data from the one of the patents is further differentiated into two information sources (Patent T2 and Patent T4). The differentiation was necessitated by additional changes to the post-processing parameters that were implemented.
@@
Name of Data Source Number of Data Samples Source of Data
Patent T2 $20$ Table 2 from Fujiwara, Teruhiko, Hiroaki Machida, and Hideyuki Yoshikawa. "Rare earth-cobalt permanent magnet." U.S. Patent Application No. 14/643,875. [US20150262740A1]
Patent T4 $21$ Table 4 from Fujiwara, Teruhiko, Hiroaki Machida, and Hideyuki Yoshikawa. "Rare earth-cobalt permanent magnet." U.S. Patent Application No. 14/643,875. [US20150262740A1]
Prop of SmCeCoFeCu $30$ Senno, Harufumi, and Yoshio, Tawara. "Permanent-magnet properties of Sm-Ce-Co-Fe-Cu alloys with compositions between 1-5 and 2-17." IEEE Transactions on Magnetics 10.2 (1974): 313-317. [senno1974permanent]
Total $71$
@@ 5.3 Model Development
Similar to the FeCrAl model development approach, GP and LVGP models are trained on the collected data. The common inputs for the GP and LVGP models are the composition of the alloy constituents (Samarium, Zirconium, Copper, Iron, Cerium, Cobalt) defined as the weight percentages (wt%) and the testing temperature of the alloy, leading to seven (7) unique input parameters. The energy density and power capacity of the magnet are experimentally measured and quantified through energy product (EP). As with the previous case study, the inputs for the LVGP contain an additional qualitative input characterizing each source of data as distinct categorical level along all the other 7 inputs used for the GP model. The outputs for both the models remain the same.
@@ 5.4 Results and Discussion
Two separate studies are conducted as previously done with FeCrAl study. Within the first study, generalizability, and the prediction capability of the models through one source that could be controlled are investigate. In this case, although there are no in-house experiments that could be controlled, Prop of SmCeCoFeCu is selected as in-house experiments (reference information source) as it is the source with the highest number of available samples. First, generalizability of GP and LVGP is compared through 5-fold cross validation using the 65 samples out of 71 available samples, and testing the models over a 6 sample separate testing set that are selected from the Prop of SmCeCoFeCu source. For CV tasks in LVGP, stratified sampling is used to make sure datasets used for training and testing represent the overall distribution of the information sources. The training dataset and testing dataset remain identical for both models. The results of this study are highlighted in Section 5.4.1.
@@
In the second study, a separate single-source GP model is built using only the Prop of SmCeCoFeCu source and compared with the multi-source LVGP model. The performances of the models are compared according to the same 6 separate testing set. For ease of notation, this single source GP model is named as GP-Prop. The results of this study are demonstrated in Section 5.4.2.
5.4.1 Multi-Source Data Fusion Comparison: LVGP and GP
@@
Table 6 summarizes the performance of the GP and LVGP model on the SmCoFe alloy dataset. From the comparison of the NRMSE across the training and cross validation datasets, it can be observed that the multi-source LVGP model shows significant improved performance over the GP model. Therefore, it can be concluded that explicitly including the source of data could enable accurate and generalizable models. Looking at the testing NRMSE metric, it is observed once again that that LVGP model outperforms the GP model, demonstrating it’s predictive capabilities. The parity plots from the CV study are also shown in Supplementary Figure 3. Moreover, from 45-degree parity plots, it is observed that both models make similar predictions on both the testing and training set (Panels A and B of Figure 9). On the other hand, it can be clearly observed that LVGP provides lower prediction uncertainties as the predictions have lower confidence bounds. Once again, this could be highly advantageous for further downstream materials design optimization studies as the information from other sources can provide a more informed and confident ML model.
@@
Table 6: Predictive performance comparison between GP, LVGP, and GP-Prop models on SmCoFe Alloys
Model   Type Mean Training NRMSE Mean CV NRMSE Testing NRMSE
GP $0.098$ $0.285$ $0.078$
LVGP $0.099$ $0.155$ $0.062$
GP-Prop $0.258$
@@
Figure 9: Predictive performance comparison between (a) GP, (b) LVGP, and (c) GP-Prop models. First row shows the parity plots of the models with respect to their respective training and Prop SmCeCoFeCu testing sets. The x and y axis show the true and predicted energy potential (EP), respectively, and error bars demonstrate the prediction uncertainty associated with each prediction. Second, third and fourth row show the Mean, Standard Deviation and 3D Response Surfaces of the models at 23%<Sm<26%, 55%<Co<63%, 1.0%<Zr<2.0%, Cu = 8.0%, Fe = 19.0%, Hf = 0%, Ce = 0% and Temperature = 25C, respectively
@@
Extending the investigation of the models, the response surface from both the models are plotted across the same input space of the Prop SmCeCoFeCu source through Panels A-B in Figure 9. The response surfaces (with Prop SmCeCoFeCu as reference due to details and nuances provided in the source) of the models are smooth and well-behaved. The well-behaved nature of the response surface can be attributed to the low dissimilarity measure of each source given through the latent space in Figure 10. This could indicate that all the sources of information possess high similarities. It is highlighted that since the experiments for Patent T2 and Patent T4 have been provided through the same Patent [US20150262740A1], it is plausible to assume that the underlying unknown physical experimental parameters are the same for both sources of information and that LVGP was able to identify this explicitly. Furthermore, in addition to providing better predictions, a case is still made for the need of multi-source data fusion through LVGP. Within the highlighted compositional space of the response surface, the location of maximum energy product from GP and LVGP are different. For the GP model, the maximum energy product is around 17.75 MGOe with a prediction uncertainty of around 2.5 MGOe, approximately at Co(wt%) = 60. For LVGP models centered at Prop SmCeCoFeCu, the maximum energy product is around 18.00 MGOe with an uncertainty of 0.50 MGOe at approximately Co(wt%) = 62. This nuanced difference is seemingly tied to the source of the data as datasets and models are identical except for inclusion of data source. In addition to higher mean predictions provided by the LVGP compared to the GP model, the predictions also come with much lower uncertainty. Furthermore, for the cases where the designer has access and control over one source of information for further design optimization, the optimum candidates provided by the GP model can be misleading and can misguide the researchers towards wrong optimum designs since there is no way to explicitly differentiate the information sources. Therefore, the reference source selection capability present in the LVGP plays an essential role in plausibly guaranteeing the optimized behavior as it is tied to the most informative source of data (identical procedures can be followed). Additionally, within the compositional space there is a difference of 2% in Co (wt%) for the maximum EP property between both models. This shift in the composition of a key rare earth element can have a significant impact in realizing such a superior alloy and downstream cost and supply chain setup.
@@
Figure 10: The latent space of SmCoFe information sources from the LVGP model. Prop SmCeCoFeCu is denoted as the reference source (*) for dissimilarity metric, $D$, calculations
5.4.2 Comparison of the Multi-Source LVGP with the Single-Source GP-Prop
@@
To overcome the challenge of performing optimization in the unknown information source space, one can argue to build GP models on each information source individually. To demonstrate the effectiveness of multi-source LVGP once again, it is compared against a single-source GP model that is built on the Prop SmCeCoFeCu source only. The single-source GP model, which is previously denoted as GP-Prop, uses the entire available training data from the Prop SmCeCoFeCu source (24 samples) for model building, and later tested on the separate 6 samples obtained from the Prop SmCeCoFeCu once again. The prediction metrics and the parity plots on testing data are shown in Table 6 (Model Types: LVGP and GP-Prop) and Panels B and C of Figure 9, respectively. From the NRMSE and parity plot results, a clear advantage of using LVGP is observed as it provides much better predictions with lower prediction uncertainties. Furthermore, looking at the response surfaces of both models, Panels B and C of Figure 9, a tone stark difference is seen in the uncertainty. For LVGP, the uncertainty for majority of the surface is at 0.5 MGOe. For th GP-Prop, the uncertainity for the majority of the surface is at 1.5 MGOe, three times of what is observed for LVGP. The high reduction in uncertainty of LVGP model comes from adding in other data-sources. This phenomenon could be a crucial differentiator when designing next generation materials with restrictions on data-availability driven by manufacturing and experimental testing capabilities.